VIBRONIC AND ROTATIONAL STRUCTURE OF THE ASYMMETRIC C-H STRETCH OF METHOXY AND d_{2}-METHOXY RADICALS.

XIAOYONG LIU, \underline{VADIM L. STAKHURSKY}, ERIC D. OLMON, VLADIMIR A. LOZOVSKY, TERRY A. MILLER, C. BRADLEY MOORE, Department of Chemistry, The Ohio State University, 120 W. 18th Avenue, Columbus, OH 43210.

Multimode Jahn-Teller analysis of the e modes in the X̃ state of the CH_{3}O radical has been initiated using the SOCJT program^{a} with special attention being paid to the asymmetric C-H stretch mode (n_{4}) for which there are several relevant experimental observations. Dispersed Fluorescence (DF) spectra have been recorded via excitation of Ã ^{2}A_{1} 3^{n}4^{1} (n = 1, 2, 3) combination levels for CH_{3}O. DF spectra have also been obtained upon 3^{n}(4′)^{1}(n = 1, 2) LIF excitation for the CHD_{2}O isotopomer. For both isotopomers the observed spectra reveal the quartet vibronic structure in the C-H stretch mode region. The observed quartet structure exhibits a large (∼130 cm^{−1} for CH_{3}O and ∼160 cm^{−1} for CHD_{2}O) separation; each component is further split into doublets (∼25 cm^{−1} and ∼35 cm^{−1} respectively for CH_{3}O and CHD_{2}O). In addition, the rotationally resolved structure of the ν_{4} C-H stretch band of CH_{3}O, recorded via stimulated emission pumping technique was compared with that recently obtained from IR spectra^{b}. The implications of the rovibronic structure for the vibronic assignments will be discussed.

aT. A. Barckholtz, T. A. Miller, “Quantitative insights about molecules exhibiting Jahn-Teller and related effects”, Int. Rev.

Phys. Chem., 17, 435, 1998

bJia-xiang Han, Yu. G. Utkin, Hong-bing Chen, L. A. Burns, R. F. Curl “High-resolution infrared spectra of the C-H asymmetric stretch vibration of jet-cooled methoxy radical (CH_{3}O)”, J. Chem. Phys., 14, 6538, 2002